The starting and ending balances are $200 and $300 at Year 1 and Year 6, respectively. What is the CAGR(%) ?
Answer #1

Ending balance (SV)- $200 , End of Year 1

Ending balance - $300, End of Year 6

Total periods, N = 5

CAGR = ( EV / SV )^1/N - 1

= (300/200) ^0.2 - 1

= 1.08447-1 = 0.08447 or 8.447%